There are currently 7,456 police officers in a particular state. 3,211 of them are females. What percent of the officers are female? What proportion of the officers are female? What percent are males? What proportion are males?
Answer by josgarithmetic(39620)

                   HOW MANY               FRACTION            PERCENT

Male Officers      7456-3211=4245        4245/7456=            56.934

Female Officers     3211                 3211/7456=            43.07

All Officers        7456                    1                 100

Percent means, "parts in one hundred".
Multiply the fraction by 100, and that is the percentage.